Ellen electron y protus proton are traveling through space in a straight line at a distance of 1cm from each other, when suddenly it gets in contact with an magnetic field with a magnetic indudction of 1T. If we assume that they are equally charged and perpendicular to the field, what is the velocity of the particles?
1
Expert's answer
2019-05-21T08:50:27-0400

The magnetic force supplies the centripetal force


"Bqv=\frac {mv^2}{r} (1)"

where r is the circle radius

Using (1) we get


"v=\frac { Bqr} {m} (2)"

Using (2) we have


for electron: v=1.8×1011 r,

for proton: v=9.6×107 r